On January 1, 2021, the Marjlee Company began construction of an office building to be used as its corporate headquarters. The building was completed early in 2022. Construction expenditures for 2021, which were incurred evenly throughout the year, totaled $7,200,000. Marjlee had the following debt obligations which were outstanding during all of 2021:


Construction loan, 10%$1,800,000
Long-term note, 9%
2,400,000
Long-term note, 6%
4,800,000


Required:
Calculate the amount of interest capitalized in 2021 for the building using the specific interest method.
Answer #1

Answer:

Step 1: Figure out your average cost of borrowing = (24,00,000*0.09 + 48,00,000*0.06)/72,00,000 = 7%
Step 2: Figure out average expenditure during year = $72,00,000/2 = $36,00,000
Step 3: Capitalized interest first eats up the specific construction loan so you had 10%*18,00,000 = $180,000
Then the next 18,00,000 is at the average borrowing rate = 7%*18,00,000 = $126,000
Total capitalized interest = $306,000
Answer #2

180,000+ 126,000= 306,000

answered by: anonymous